Find the projection of the line segment joining the points (1, βˆ’1, 6) and (4, 3, 2) on the line (π‘₯βˆ’4)/3 = βˆ’π‘¦ = 𝑧/5 .


1
Expert's answer
2022-02-14T17:08:30-0500

The projection of the segment joining the points "P(x_1, y_1, z_1)" and "Q(x_2, y_2, z_2)" on a line having direction cosines "l, m, n" is "|l(x_2 \u2013 x_1) + m(y_2 \u2013 y1) + n(z_2 \u2013 z_1)|."


"\sqrt{(3)^2+(-1)^2+(5)^2}=\sqrt{35}"

The projection of the given segment is


"|\dfrac{3}{\sqrt{35}}(4-1)-\dfrac{1}{\sqrt{35}}(3+1)+\dfrac{5}{\sqrt{35}}(2-6)|"

"=\dfrac{15}{\sqrt{35}}"